The supplement of an angle is 12 more than 3 times its compliment. The angle measure is ....?
Answer by TimothyLamb(4379)

---
x = angle
---
180 - x = 3(90 - x) + 12
180 - x = 270 - 3x + 12
2x = 102
---
answer:
x = 51 degrees
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php